Reports 1-1 of 1 Clear search Modify search
MIF (General)
hirotaka.yuzurihara - 13:53 Thursday 22 May 2025 (33869) Print this report
Initial alignment 20250522 and the interferometer reached the PRFPMI_RF_LOCK

[Kenta, Yokozawa, Yuzurihara]
We performed the initial alignment. During taking the alignment for the PRMI, the earthquake happened and the BS was fluctuating ~3urad. We wait for a while and finished the initial alignment at 13:37.
We requested the PRFPMI_RF_LOCK to the LSC_LOCK guardian and confirmed that the interferometer can reach the PRFPMI_RF_LOCK at 13:50.
One concern is that the IR TRANS powers seem to be low <110.

Images attached to this report
Comments to this report:
kenta.tanaka - 18:25 Thursday 22 May 2025 (33873) Print this report

We suceeded in restoring PRFPMI DC lock with 10W after the ITMX release by warming up.

Fig. 1, 2, 3, and 4 show the OLTFs of DARM, MICH, PRCL, and CARM, respectively. DARM, MICH and CARM seem to be lower to at most ~1dB than previous ones.  PRCL seems to be almost consistent with the previous one.

Therefore, LSCs seems to be fine.

Images attached to this comment
Search Help
×

Warning

×